# A Multi-Messenger Search for the Supermassive Black Hole Binary in 3C 66B with the Parkes Pulsar Timing Array

### v1

@@ 4.2 Limits on the SMBHB chirp mass at 3C 66B
We also report the result of our reanalysis of the 3-mm data from Iguchi et al. (2010). Following the methodology in Section 3.4, we obtain a posterior on the log-10 chirp mass of the SMBHB at 3C 66B. It is shown as the black shaded area in Figure 2. The posterior is skewed towards higher masses, consistently with the uncertainties reported in Iguchi et al. (2010) and presented in Table 1.
Figure 2: Observational constraints on the chirp mass $\mathcal{M}$ of an SMBHB at 3C 66B. The histograms show a posterior for $\mathcal{M}$ based on the analysis of PPTA DR3. The dashed lines show the respective upper limits at $95\%$ credibility based on the astrophysically-motivated prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,9.5)~[\log_{10}M_{\odot}]$. The dotted lines show the limits based on the conservative prior $\pi(\mathcal{M})=\mathcal{U}(10^{7.5},10^{9.5})~[M_{\odot}]$. In green, we show the results of the analysis with only the Earth term of the signal modeled. In blue, we show the results based on modeling both the Earth term and the pulsar term of the CGW. The black solid line with arbitrary density normalization represents the posterior we obtain from the 3-mm data analyzed in Iguchi et al. (2010). The vertical orange line and the band correspond to the value and the uncertainty reported in Iguchi et al. (2010).
4.3 Limits on the gravitational wave strain from 3C 66B
Instead of calculating the strain amplitude $h_{0}$ of the CGW from 3C 66B based on the luminosity distance using Equation 2, we now treat it as a free parameter. This way, the distance to 3C 66B is not involved in the calculation. Similar to the chirp mass prior, we use a prior $\pi(\log_{10}h_{0})=\mathcal{U}(-17.0,-12.5)$, which is more consistent with the decreasing probability of an SMBHB with a high $h_{0}$. The other priors are as in Section 4.2. The posterior on $h_{0}$ is shown in Figure 3. With only the Earth term, we find an upper limit at 95% credibility to be $\log_{10}h_{0}<-14.12$. When including the pulsar term, we find a more constraining limit, $\log_{10}h_{0}<-14.44$. Using the more conservative prior, $\pi(h_{0})=\mathcal{U}(10^{-17.0},10^{-12.5})$, we find an upper limit for the Earth term of $\log_{10}h_{0}<-13.86$ and an upper limit of $\log_{10}h_{0}<-14.20$ when including the pulsar term. As in Section 4.2, the range of $h_{0}$ of the EM model from Iguchi et al. (2010) is partially ruled out in all cases.
It is also of interest to compare the strain amplitude of a CGW signal from 3C 66B with the expectation for the GWB at the EM-predicted frequency of the binary. Goncharov et al. (2024) proposes a fiducial prior on the characteristic strain amplitude of the background, $h_{\text{c}}(f)=A(f~\text{yr})^{-2/3}$, characteristic of SMBHBs. The prior is based on black hole masses inferred from kinematic observations and galaxy velocity dispersion. Based on this prior, we calculate $h_{0,\text{GWB}}=2h_{\text{c}}/\sqrt{32/5}$, to revert the root-mean-square averaging over SMBHB inclination and polarization assumed for the background (see Equation 8 in Burke-Spolaor et al. (2019)). The best-fit $h_{\text{c}}(f=\text{yr}^{-1})$ of the GWB from the data of the European Pulsar Timing Array by Goncharov et al. (2024) is within $3\sigma$ interval of the prior, above the prior mean. Thus, the prior is consistent with observations.
  • The EM model of 3C 66B corresponds to a loud source in the expected GWB.
  • A fraction of the parameter space of the EM model for 3C 66B, which we have ruled out, is also not supported by the predicted GWB amplitude.
  • Our limits at $f_{\text{CGW}}$ of 3C 66B are, on average, higher than the expected GWB because we are limited by the white noise at the frequency of 3C 66B.
The value $h_{0,\text{GWB}}$ can also be interpreted as the equivalent expected CGW amplitude of the background if it is dominated by one source (3C 66B) at 3C 66B CGW frequency. This scenario is possible, given the scaling of the number of SMBHB sources contributing to the background with frequency. This may also lead to a deviation from the power-law model for $h_{\text{c}}(f)$ at the frequency of 3C 66B, depending on the number density of SMBHBs. It is also worth noting that Zhu et al. (2019) find that the EM model for 3C 66B is in tension with the observed GWB amplitude based on the expected SMBHB merger rate.
Figure 3: Observational constraints on the CGW strain amplitude, $h_{0}$, of an SMBHB at 3C 66B. The histograms show a posterior for $h_{0}$ based on the analysis of PPTA DR3. The dashed lines show the respective upper limits at $95\%$ credibility based on the astrophysically-motivated prior $\pi(\log_{10}h_{0})=\mathcal{U}(-17.0,-12.5)$. The dotted lines show the limits based on the conservative prior $\pi(h_{0})=\mathcal{U}(10^{-17.0},10^{-12.5})$. In green, we show the results of the analysis with only the Earth term of the signal modeled. In blue, we show the results based on modeling both the Earth term and the pulsar term of the CGW. The black solid line with arbitrary density normalization represents the theoretically expected gravitational wave background contribution at the frequency of 3C 66B. The vertical orange line and the band correspond to the value and the uncertainty reported in Iguchi et al. (2010).
4.4 Upper limits as a function of CGW frequency
Following Arzoumanian et al. (2020b), for a broader perspective on how PPTA DR3 limits on $\mathcal{M}$ depend on $f_{\text{GW}}$ for any SMBHB at the position of 3C 66B, we show Figure 4. For this calculation, we use the same prior as in Section 4.2, except the prior on CGW frequency is now $\pi(\log_{10}f_{\text{GW}})=\mathcal{U}(-9.0,-7.0)~[\log_{10}\text{Hz}]$. We also only model the Earth term of the CGW. The figure highlights that the PPTA is well-suited to target 3C 66B, as the value suggested by the EM model is at the frequency to which our targeted search is most sensitive. Consistent with Figure 2, when modeling the pulsar term and using astrophysically motivated priors, we are close to ruling out the best-fit value of the EM model. We observe the highest density of posterior samples at around $f_{\text{GW}}=\text{yr}^{-1}$, which corresponds to uncertainties associated with measuring pulsar positions. There is also a minor over-density of posterior samples at $f_{\text{GW}}=(2~\text{yr})^{-1}$, which corresponds to uncertainties associated with measuring pulsar parallax. Given that the posterior is not uniform at these frequencies, some of the uncertainties may be systematic.
Figure 4: Posterior for the chirp mass $\mathcal{M}$ and frequency $f_{\text{GW}}$ of 3C 66B. The color map shows the number of posterior samples in the $\mathcal{M}$-$f_{\text{GW}}$ space. The best-fit value from Iguchi et al. (2010) is shown as a red star. The blue solid line shows the respective upper limit at $95\%$ credibility based on the astrophysically-motivated prior $\pi(\log_{10}\mathcal{M})=\mathcal{U}(7.5,10.5)~[\log_{10}M_{\odot}]$. The yellow dashed line shows the limits based on the conservative prior $\pi(\mathcal{M})=\mathcal{U}(10^{7.5},10^{10.5})~[M_{\odot}]$.
4.5 Frequentist Search
To confirm the non-detection of a CGW, we perform the frequentist targeted search based on the $\mathcal{F}$ statistic (Ellis et al., 2012), a frequentist measure of evidence for a CGW in our data. In particular, we chose to calculate the $\mathcal{F}_{\text{e}}$ statistic, which is based on only modeling the Earth term of CGW. This statistic is a function of CGW frequency and sky position, which we set to the 3C 66B EM model values. We find the value of $\mathcal{F}_{\text{e}}=4.90$, corresponding to the false alarm probability $\text{FAP}=0.50$. Ellis et al. (2012) suggest a threshold for FAP to be $10^{-4}$ to claim a detection. The trials factor for our search is $1$. The background distribution of the $\mathcal{F}_{\text{e}}$ statistic is the chi-squared distribution with $4$ degrees of freedom and is displayed in Figure 5 along with the value of the $2\mathcal{F}_{\text{e}}$ statistic. Based on this null distribution, we find a $p\text{-value}=0.04$ which corresponds to about $2.1\sigma$. Thus, the $\mathcal{F}_{\text{e}}$ statistic value we find is consistent with noise.
Figure 5: The background distribution of the $\mathcal{F}_{\text{e}}$ statistic, which is a chi-squared distribution with $4$ degrees of freedom. The green vertical dashed line is the $2\mathcal{F}_{\text{e}}$ statistic calculated at the frequency and sky position of 3C 66B.
4.6 Cosmology based on targeted searches with PTAs
Standard sirens are compact binaries that emit gravitational and electromagnetic (EM) radiation simultaneously, allowing for measurement of the universe’s expansion rate (Schutz, 1986). Redshift of the host galaxy of a standard siren from EM information enables the determination of its recessional velocity $v$. Whereas the proper distance $D$ is measured based on the gravitational wave signal. The Hubble constant, which quantifies the universe’s expansion at the present time, is
$H_{0}=\frac{\nu}{D}.$ (6)
The primary challenge in the standard siren approach is localizing the electromagnetic (EM) counterpart of a gravitational wave. To this date, the first observed binary neutron star inspiral, GW170817 (Abbott et al., 2017a), is the only gravitational wave signal for which an EM counterpart is established: galaxy NGC 4993. The sky position uncertainty for GW170817 is 28 $\text{deg}^{2}$ at 90% credibility. For PTAs, the sky position uncertainty for all-sky CGW searches is expected to be less precise. Moreover, the signals are not transient, which makes identification of the host galaxy even more difficult. For example, IPTA-DR3 simulations (116 pulsars, 20-year baseline) from Petrov et al. (2024) found that for an ideal case where the signal-to-noise ratio (SNR) is of $\text{SNR}=15$, 90% credible area would range from 29 $\text{deg}^{2}$ to 241 $\text{deg}^{2}$. However, for a realistic case, with $\text{SNR}=8$, the 90% credible area increases from 287 $\text{deg}^{2}$ to 530 $\text{deg}^{2}$. Petrov et al. (2024) concludes that without better CGW localization, host galaxy searches are impractical. Truant et al. (2025) have additionally simulated a 30-year SKA-PTA observation and found that out of the $\approx 35$ CGW detected, only 35% have a 90% credible area <100<100 $\text{deg}^{2}$. They conclude that observations with an SNR >25>25 will have fewer than 100 host galaxy candidates. Even in simulations which look at a very high SNR, such as those performed by Zhu et al. (2016), with a 15-year baseline and 30 pulsars, a strong signal with $\text{SNR}=100$ yields a localization of $\sim 40$ $\text{deg}^{2}$. Therefore, standard siren cosmology based on all-sky CGW searches is not feasible with currently operating PTAs.
@@ 4.6 Cosmology based on targeted searches with PTAs
Targeted PTA searches for SMBHB candidates eliminate the problem of finding the electromagnetic counterpart of a gravitational wave signal. For galaxy 3C 66B, an SMBHB candidate source, we know the sky position and redshift $z$ to a very high precision (Hunt et al., 2021; van den Bosch et al., 2015). Because $z=0.02\ll 0.1$ and because we do not expect stringent constraints on $D$, we assume $D=D_{\text{L}}$ and use the linear approximation $v=cz$ for simplicity. To develop high-precision cosmology with PTAs as part of future work, it is necessary to use full expressions for $D$ and $v$, while also correcting the recession velocities based on peculiar velocities of galaxies (Abbott et al., 2017b). In any case, the key requirement of a PTA in cosmology, based on targeted CW searches, is to constrain the luminosity distance of an SMBHB candidate.
@@
Therefore, to obtain a posterior on $H_{0}$, we place constraints on the luminosity distance $D_{\text{L}}$ to 3C 66B with the PPTA DR3. We impose a prior $\pi(D_{\text{L}})=\mathcal{U}(10^{-2},10^{3})$ Mpc. We chose a prior on $\log_{10}f_{\text{GW}}$ as in Section 4.2, and we chose a prior on $\log_{10}\mathcal{M}$ as in Section 4.1. Based on the posterior on $\pi(D_{\text{L}})$ and Equation 6, we obtain a posterior on $H_{0}$ in Figure 6. The values $H_{0}$ are shown along the bottom horizontal axis. The values of $\pi(D_{\text{L}})$ corresponding to $H_{0}$ based on Equation 6 and the known redshift of 3C 66B are shown along the top horizontal axis. Because we do not detect a gravitational wave signal, we rule out the nearest luminosity distances at which the SMBHB candidate is located. Our lower limits on $D_{\text{L}}$ determine our upper limits on $H_{0}$. Our choice of $\pi(D_{\text{L}})$ is presented as the toy model, which is why we do not currently report the $95\%$ credible upper-limit value for $H_{0}$. Overall, we find that the PTA constraints on $H_{0}$ based on our targeted search for the SMBHB at 3C 66B are not competitive with those of the state-of-the-art cosmological experiments. However, a detection of CGWs from an SMBHB in the targeted search, especially with the resolved pulsar term, may prove to be useful.
@@
Figure 6: Posterior on the Hubble constant. The prior is shown as the blue histogram and the posterior as the orange histogram. The uncertainty regions for the Hubble constant from Planck Collaboration et al. (2020) and the SH0ES collaboration (Riess et al., 2022) are plotted as the purple and green regions, respectively. The luminosity distance of 10 Mpc is also plotted for reference as the black vertical dashed line, showing that the nearest distances from the Earth are ruled out.
5 Conclusion
@@
Using the PPTA DR3, we conduct a targeted multi-messenger search for an SMBHB at the center of 3C 66B. We search for a CGW signal by performing model selection, where we compare a model with a CGW signal to the null hypothesis. With priors carefully matched to the EM model, we find a log Bayes factor of $\ln\mathcal{B}=0.22$. This suggests we can neither confirm nor rule out the EM model. We also find frequentist $\mathcal{F}_{e}=5.04$ with an associated false alarm probability of $\text{FAP}=0.50$, which confirms non-detection of the signal. Based on the background estimation, we find that $\mathcal{F}_{e}=5.04$ corresponds to the p-value $p=0.04$, about $2.1\sigma$.
@@
Following non-detection, we place upper limits on the mass as $\mathcal{M}<6.05\times 10^{8}~M_{\odot}$ and strain amplitude $\log_{10}(h_{0})<-14.44$, at 95% credibility. For both the mass and the strain, we rule out about 50% of the parameter space of the EM model. Using conservative priors, we still rule out a portion of the EM model, finding a $95\%$ upper limit on the chirp mass of $\mathcal{M}<9.55\times 10^{8}~M_{\odot}$ and for the strain amplitude we find a value of $h_{0}<-14.20$. Finally, we demonstrate a proof of principle that it is possible to use firmly established SMBHB candidates in targeted searches with PTAs as standard sirens, complementary to high-precision cosmology. In particular, we calculate the posterior of the luminosity distance to 3C 66B, based on which we obtain constraints on the Hubble constant.
6 Acknowledgments
@@
Appendix A A continuous gravitational wave from an SMBHB
Let us start by considering a plane wave traveling from an SMBHB to the Solar System barycenter (SSB) to which pulse arrival times are referenced. A unit vector that points from the CGW source to the SSB is
@@ Appendix A A continuous gravitational wave from an SMBHB
The wave is characterized by the metric strain rank-$2$ tensor in the transverse traceless gauge, indexed with $(a,b)$,
@@
Polarization angle $\psi$ is the angle between the natural polarization basis (at source) and the reference polarization basis (at SSB). With this, polarization tensors are expressed as
@@
The response of a PTA to the plane wave is described by the antenna pattern function (Romano & Cornish, 2017)
@@
where $\hat{u}$ is the unit vector pointing from the SSB to the pulsar. This function describes a pulsar’s geometrical sensitivity to the CGW source (Taylor et al., 2016). Pulsar timing delays $\delta t$ are then expressed as
@@
where $t$ is the time when the CGW passes the SSB and $t_{\text{p}}$ is the time when the CGW passes the pulsar. One can then geometrically relate $t$ and $t_{\text{p}}$,
@@
where $L$ as the distance to the pulsar.
For an SMBHB, at zeroth post-Newtonian order, we arrive at our model of the CGW projected onto a PTA detector,
@@
where $\iota$ is the SMBHB inclination angle, $D_{\text{L}}$ is the luminosity distance, $\omega(t)=\pi f_{\text{GW}}$ is the angular orbital frequency [^2] The SMBHB angular orbital frequency is half the CGW angular frequency $2\pi f_{\text{GW}}$., $\Phi(t)$ is the phase,
@@
is the chirp mass, $m_{1,2}$ are SMBHB component masses. The last observation in the PPTA dataset (March 8, 2022, or MJD 59646) is used as a reference time. The orbital phase and frequency of the SMBHB are then defined as,
@@
with $\Phi_{0}$ being the initial orbital phase and $\omega_{0}$ the initial orbital frequency, one may notice that it is equivalent to Equation 3.
Chirp mass and frequency are redshifted for CGW sources at cosmological distances, such that the observed “o” values are related to intrinsic source “s” values as
@@
$f_{\text{o}}$ $=\frac{f_{\text{s}}}{1+z}.$
For 3C 66B, $z=0.02\ll 1$, so given our measurement uncertainty we assume $\mathcal{M}=\mathcal{M}_{\text{o}}=\mathcal{M}_{\text{s}}$ and $f=f_{\text{o}}=f_{\text{s}}$.
Appendix B Analysis of the 3-mm data
Although key details of the analysis of the 3-mm data are provided in Iguchi et al. (2010), here we provide explicit equations that we used in our reanalysis. Throughout, we switch between the velocity $v$ and the boost parameter $\beta=v/c$, which may also have subscripts “bl,0” and “app” for the absolute binary component velocity around the center of gravity and the apparent relativistic jet velocity, respectively. First, we point out that Equations 1 and 2 from Iguchi et al. (2010) can be expressed directly in terms of $(\beta,\iota)$,

### v2

@@ V.2 Feasibility of Standard Siren Cosmology using Targeted Searches with Pulsar Timing Arrays
We propose that by taking advantage of targeted searches for SMBHB candidates, we may be able to revisit the prospect of standard siren cosmology with PTAs.
Current standard siren methods mostly focus on follow-ups from all-sky searches (Jin et al., 2025), which do not offer a priori knowledge of the source position. In contrast, targeted searches based on EM candidates assume the known sky location of the host galaxy. It is important to note, however, that targeted searches do not localize a source within a large uncertainty region, rather they test if the CGW model is consistent with the location of the candidate. This means that source misspecification is possible if there exist multiple hosts that overlap in parameter space. Further studies to determine the extent of the host identification problem and possible solutions are ongoing.
@@
As a demonstration, let us consider a toy example of obtaining constraints on the Hubble–Lemaître constant, provided precise knowledge of the sky position and redshift (Hunt et al., 2021; van den Bosch et al., 2015) and the (lack of) GW emission from 3C 66B. Because $z=0.02\ll 0.1$ and because we do not expect stringent constraints on $D$, we assume $D=D_{\text{L}}$ and use the linear approximation $v=cz$ for simplicity. To develop high-precision cosmology with PTAs as part of future work, it is necessary to use full expressions for $D$ and $v$, while also correcting the recession velocities based on peculiar velocities of galaxies (Abbott et al., 2017b). In any case, the key requirement of a PTA in cosmology, based on targeted CGW searches, is to constrain the luminosity distance of an SMBHB candidate.
@@
Therefore, to obtain a posterior on $H_{0}$, we place constraints on the luminosity distance $D_{\text{L}}$ to 3C 66B with the PPTA DR3 data. We impose a prior $\pi(D_{\text{L}})=\mathcal{U}(10^{-2},10^{3})$ Mpc. We chose a prior on $\log_{10}f_{\text{GW}}$ as in Section IV.2, and we chose a prior on $\log_{10}\mathcal{M}$ as in Section IV.1. Based on the posterior on $\pi(D_{\text{L}})$ and Equation 6, we obtain a posterior on $H_{0}$ in Figure 6. The values $H_{0}$ are shown along the bottom horizontal axis. The values of $\pi(D_{\text{L}})$ corresponding to $H_{0}$ based on Equation 6 and the known redshift of 3C 66B are shown along the top horizontal axis. Assuming that the EM model is correct, since we do not detect a GW signal, we rule out the nearest luminosity distances. This, in turn, limits the highest values of $H_{0}$. We find that the PTA constraints on $H_{0}$ based on the targeted search generally recover the prior, other than the largest values of $H_{0}$. Our choice of $\pi(D_{\text{L}})$ is presented as a toy model, which is why we do not currently report the $95\%$ credible upper-limit value for $H_{0}$. Because we do not claim any detection of a CGW signal from 3C 66B, we emphasize that these constraints are subject to the validity of the CGW model, where 3C 66B is a true SMBHB. In other words, the EM model for 3C 66B must be correct in order for the constraints to be meaningful. Nevertheless, a strong detection of a CGW from an SMBHB in a targeted search, especially with the resolved pulsar term, may prove to be useful in constraining $H_{0}$, and investigations into this matter are underway.
@@
Figure 6: Posterior on the Hubble–Lemaître constant. The prior is shown as the blue histogram and the posterior as the orange histogram. The uncertainty regions for the Hubble–Lemaître constant from Planck Collaboration et al. (2020) and the SH0ES collaboration (Riess et al., 2022) are plotted as the purple and green regions, respectively. A luminosity distance of 10 Mpc is also plotted for reference as the black vertical dashed line.
VI Conclusion
@@
Using PPTA DR3, we conduct a targeted multimessenger search for an SMBHB at the center of 3C 66B. We search for a CGW signal by performing model selection, where we compare a model with a CGW signal to the null hypothesis. With priors carefully matched to the EM model, we find a log Bayes factor of $\ln\mathcal{B}=-0.0027(7)$. This suggests we can neither confirm nor rule out the EM model. We also find a frequentist statistic, $\mathcal{F}_{e}=5.04$, with an associated $\text{FAP}=0.50$, which confirms nondetection of the signal. Based on the background estimation, we find that the value of $\mathcal{F}_{e}=5.04$ corresponds to $p=0.04$, about $2.1\sigma$.
Following nondetection, we place upper limits on the mass as $\mathcal{M}<6.90\times 10^{8}~M_{\odot}$ and strain amplitude $\log_{10}(h_{0})<-14.44$, at 95% credibility.
@@
For both the mass and the strain, we rule out over 50% of the parameter space of the EM model. Using conservative priors, we still rule out a portion of the EM model, finding a $95\%$ upper limit on the chirp mass of $\mathcal{M}<9.55\times 10^{8}~M_{\odot}$ and for the strain amplitude we find a value of $h_{0}<-14.20$. Finally, we demonstrate a new methodology for a joint-likelihood approach for the analysis of 3C 66B as well as a proof of principle, showing that it may be possible to use firmly established SMBHB candidates in targeted searches with PTAs as standard sirens, complementary to high-precision cosmology.
VII Acknowledgments
@@
Appendix A A Continuous Gravitational Wave from a Supermassive Black Hole Binary
Let us start by considering a plane wave traveling from an SMBHB to the SSB to which the pulse arrival times are referenced. A unit vector that points from the CGW source to the SSB is
@@ Appendix A A Continuous Gravitational Wave from a Supermassive Black Hole Binary
The wave is characterized by the metric strain rank-$2$ tensor in the transverse traceless gauge, indexed with $(a,b)$:
@@
The polarization angle $\psi$ is the angle between the natural polarization basis (at source) and the reference polarization basis (at SSB). With this, polarization tensors are expressed as
@@
The response of a PTA to the plane wave is described by the antenna pattern function (Romano and Cornish, 2017)
@@
where $\hat{u}$ is the unit vector pointing from the SSB to the pulsar. This function describes the pulsar’s geometrical sensitivity to the CGW source (Taylor et al., 2016). Pulsar timing delays $\delta t$ are then expressed as
@@
where $t$ is the time when the CGW passes the SSB and $t_{\text{p}}$ is the time when the CGW passes the pulsar. One can then geometrically relate $t$ and $t_{\text{p}}$, as
@@
where $L$ is the distance to the pulsar.
For an SMBHB, at zeroth post-Newtonian order, and choosing a coordinate system where $\psi=0$, we arrive at our model of the CGW projected onto the PTA detector:
@@
where $\iota$ is the SMBHB inclination angle, $D_{\text{L}}$ is the luminosity distance, $\omega(t)=\pi f_{\text{GW}}$ is the angular orbital frequency [^2] The SMBHB angular orbital frequency is half the CGW angular frequency $2\pi f_{\text{GW}}$., $\Phi(t)$ is the phase, and
@@
is the chirp mass, where $m_{1,2}$ are the SMBHB component masses. The last observation in the PPTA dataset (2022 March 8, or MJD 59646) is used as the reference time. The orbital phase and frequency of the SMBHB are then defined as
@@
with $\Phi_{0}$ being the initial orbital phase and $\omega_{0}$ the initial orbital frequency; one may notice that this equation is equivalent to Equation 3.
The chirp mass and frequency are redshifted for CGW sources at cosmological distances, such that the observed (“oo”) values are related to intrinsic source (“ss”) values as
@@
$f_{o}$ $=\frac{f_{s}}{1+z}.$
For 3C 66B, $z=0.02\ll 1$, so given our measurement uncertainty we assume $\mathcal{M}=\mathcal{M}_{o}=\mathcal{M}_{s}$ and $f=f_{o}=f_{s}$.
Appendix B Analysis of the 3 mm Data
Although key details of the analysis of the 3 mm data are provided in Iguchi et al. (2010), here we provide explicit equations that we used in our reanalysis. Throughout, we switch between the velocity $v$ and the boost parameter $\beta=v/c$, which may also have subscripts “bl,0” and “app” for the absolute binary component velocity around the center of gravity and the apparent relativistic jet velocity, respectively. First, we point out that Equations 1 and 2 from Iguchi et al. (2010) can be expressed directly in terms of $(\beta,\iota)$, as